# Conformal prediction for causal effects of continuous treatments

### v2

@@ 4.1 Scenario 1: Known propensity score
Lemma 1 ((Gibbs et al., 2023)).
Let $\mathcal{F}$ define a finite-dimensional function class that includes the function $f$ characterizing the shift in the (potentially unknown) propensity function $\pi$ (see Eq. equation 2). Define the distribution-shift-calibrated $(1-\alpha)$-quantile of the non-conformity scores as
$\hat{g}_{S}(X_{n+1}):=\operatorname*{arg\,min}\limits_{g\in\mathcal{F}}\frac{1}{n-m}\left(\sum_{i=m+1}^{n}l_{\alpha}(g(X_{i}),S_{i})+l_{\alpha}(g(X_{n+1}),S)\right)$ (5)
@@
Building upon Lemma 1, we derive our first main result in Theorem 1. We define the finite-dimensional function class of interest as :=$\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$. It is easy to verify that all $f\in\mathcal{F}$ represent the desired propensity shift to $\tilde{\pi}(a\mid x)=\pi(a+\Delta_{A}\mid x)$ as defined in Eq. equation 2.
Theorem 1 (Conformal prediction intervals for known baseline policy).
@@
s.t. $\qquad\qquad\qquad\qquad-\alpha\leq\eta_{i}\leq 1-\alpha,\quad\forall i=m+1,\ldots,n+1,$
@@
We provide a full proof in Supplement C.2. Here, we briefly outline the underlying idea of the proof. First, we show that the function class $\mathcal{F}$ indeed satisfies Eq. equation 2 for the intervention $A^{*}(X)=A+\Delta_{A}$, and we then rewrite Eq. equation 5 as a convex optimization problem. Next, we exploit the strong duality property, we optimize over the corresponding dual problem to receive a dual prediction set with equal coverage probability. Finally, we derive $S^{\ast}$ from the dual prediction set to construct $C_{n+1}$ and prove the overall coverage guarantee. ∎
@@ 4.2 Scenario 2: Unknown treatment policy
which hinders a direct adaptation of Theorem 1 due to the inherent discontinuity of the improper function. Hence, to proceed, we make the following assumption on the propensity estimator.
@@
can be represented in terms of Eq. equation 2 through functions $f\in\mathcal{F}$.
@@
See Supplement C.1. ∎
Following the motivation in scenario  1, we thus aim to estimate the $(1-\alpha)$-quantile of the non-conformity scores under the distribution shift in Lemma 1. We can reformulate this problem as
@@
s.t. $\qquad\qquad S_{i}-\frac{c_{a}}{\sqrt{2\pi}\sigma}\frac{\exp{(-\frac{(a_{i}-a^{\ast})^{2}}{2\sigma^{2}})}}{\hat{\pi}(a_{i}\mid x_{i})}-u_{i}+v_{i}=0,\quad\forall i=m+1\ldots,n+1,$
$\qquad\qquad u_{i},v_{i}\geq 0,\quad\forall i=m+1\ldots,n+1$
for the imputed score $S_{n+1}=S$. As the score is unknown, computing the CP interval would require solving equation PS for all $S\in\mathbb{R}$, yet which is computationally infeasible. As a remedy, we previously exploited properties of the dual optimization problem and the Lagrange multipliers of the convex problem in Theorem 1 to efficiently compute the CP intervals (see the proof of Theorem 1 in Supplement C). However, the present non-convex problem does not automatically allow for the same simplifications. Instead, we now present a remedy for efficient computation of the CP intervals in the following lemma. We prove Lemma 3 in Supplement C.
Lemma 3.
The problem equation PS is Type-I invex and satisfies the linear independence constraint qualification (LICQ).
Lemma 3 allows us to derive properties of the present non-convex optimization problem in terms of the Karush-Kuhn-Tucker (KKT) conditions. For this, we note that the fulfillment of the LICQ serves as a sufficient regularity condition for the KKT to hold at any (local) optimum of equation PS. Combined with the Type-I invexity of the objective function and the constraints, the KKT conditions are not only necessary but also sufficient for a global optimum. As a result, we can employ the KKT conditions at the optimal values [^3] We provide an interpretation of the optimal values in Supplement E. Therein, we further discuss the implications of the proposed kernel smoothing of $\delta_{a^{\ast}}(a)$. $\sigma^{\ast}$ and $c_{a}^{\ast}$ to derive coverage guarantees of our CP interval in a similar fashion as in Theorem 1. We thus arrive at the following theorem to provide CP intervals for the scenario with unknown propensity scores.
Theorem 2 (Conformal prediction intervals for unknown propensity scores).
@@
s.t. $\qquad\qquad S_{i}-\frac{c_{a}}{\sqrt{2\pi}\sigma}\frac{\exp{(-\frac{(a_{i}-a^{\ast})^{2}}{2\sigma^{2}})}}{\hat{\pi}(a_{i}\mid x_{i})}-u_{i}+v_{i}=0,\quad\forall i=m+1\ldots,n+1,$
$\qquad\qquad u_{i},v_{i}\geq 0,\quad\forall i=m+1\ldots,n+1.$
for an imputed unknown $S_{n+1}=S$. Let $S^{\ast}$ be defined as the maximum $S$ s.t. $v_{n+1}^{S}>0$. Then, the interval
@@
4.3 Algorithm
@@ 21      $S^{\ast}\leftarrow\frac{S_{\mathrm{up}}+S_{\mathrm{low}}}{2}$;
We now use our Theorem 2 to present an algorithm for computing CP intervals of potential outcomes from continuous treatment variables under unknown propensities in Alg. 1. We present a similar algorithm for scenario  1 with known propensities and discuss the computational complexity in Supplement B.
We make the following comments: In our algorithm, an optimization solver is used to calculate $v_{n+1}$ according to Theorem 2. The specific choice of the solver is left to the user. In our experiments in Section 5, we perform the optimization via interior point methods. Further, the overall goal of our algorithm is to find the optimal imputed non-conformity score $S^{\ast}$ such that the coverage guarantees hold. This search can be implemented through suitable iterative search algorithms.
5 Experiments
Baselines: As we have discussed above, there are no baselines that directly compute finite-sample prediction intervals for potential outcomes of continuous treatments. Therefore, we compare our method against MC dropout (Gal & Ghahramani, 2016) and deep ensemble methods (Lakshminarayanan et al., 2017). Yet, we again emphasize that MC dropout is an ad hoc method with poor approximations of the posterior, which is known to give intervals that are not faithful (Le Folgoc et al., 2021).
Implementation: All methods are implemented with $\phi$ as a multi-layer perception (MLP) and an MC dropout regularization of rate 0.1. Crucially, we use the identical MLP for both our CP method and MC dropout. Hence, all performance gains must be attributed to the coverage guarantees of our conformal method. In the MC dropout baseline, the uncertainty intervals are computed via Monte Carlo sampling. In scenario 2 with unknown propensity scores, we perform the conditional density estimation through conditional normalizing flows (Trippe & Turner, 2018). Details about our implementation and training are in Supplement E.
Performance metrics: We evaluate the methods in terms of whether the prediction intervals are faithful (e.g., Hess et al., 2024). That is, we compute whether the empirical coverage of the prediction intervals surpasses the threshold of $1-\alpha$ for different significance levels $\alpha\in\{0.05,0.1,0.2\}$.
5.1 Datasets
Synthetic datasets: We follow common practice in causal ML and evaluate our methods using synthetic datasets (e.g., Alaa et al., 2023; Jin et al., 2023). The reason is the fundamental problem of causal inference, because of which counterfactual outcomes are never observable in real-world datasets. Synthetic datasets enable us to access counterfactual outcomes and, thus, to benchmark methods in terms of whether the computed intervals are faithful.
We consider two synthetic datasets with different propensity scores and outcome functions. Dataset 1 uses a step-wise propensity function and a concave outcome function. Dataset 2 is more complex and uses a Gaussian propensity function and oscillating outcome functions. Both datasets contain a single discrete confounder, a continuous treatment, and a continuous outcome. By choosing low-dimensional datasets, we later render it possible to plot the treatment–response curves so that one can inspect the prediction intervals visually. (We later also show that our method scales to high-dimensional settings as part of the real-world dataset.) Details about the data-generating processes are in Supplement E.
Medical dataset: We demonstrate the applicability of our CP method to datasets from medicine by leveraging the MIMIC-III dataset (Johnson et al., 2016). MIMIC-III contains de-identified health records from patients admitted to critical care units at large tertiary care hospitals. Our goal is to predict patient outcomes in terms of blood pressure when treated with a different duration of mechanical ventilation. We use 8 confounders from medical practice (e.g., respiratory rate, hematocrit). Overall, we consider 14,719 patients, which we randomly split into train (60%), validation (10%), calibration (20%), and test (10%). Further details are in Supplement E.
5.2 Results for synthetic datasets
We now evaluate the faithfulness of our CP intervals through a series of experiments. On each dataset, we analyze the performance of the prediction intervals in the presence of various soft interventions $\Delta_{a}\in\{1,5,10\}$ and hard interventions $a^{\ast}\in\{5x,7x,10x\}$ for each $X=x$. We average the empirical coverage across 10 runs with random seeds. The results are in Fig. 4 (dataset 1) and Fig. 5 (dataset 2).
We make the following observations. First, the intervals of our CP method comply with the targeted significance level $\alpha$ and, therefore, are faithful. Second, both MC dropout and the deep ensemble method have, in contrast, a considerably lower coverage, implying that the intervals are not faithful. This is in line with the literature, where MC dropout is found to produce poor approximations of the posterior (Le Folgoc et al., 2021). In particular, the ensemble method is highly unfaithful. Thus, we will not consider this baseline in all of the following experiments. Third, our method has only a small variability in terms of empirical coverage, whereas the empirical coverage of MC dropout varies greatly. This corroborates the robustness of our method. Fourth, the results are consistent for both datasets. In sum, this demonstrates the effectiveness of our proposed CP method.
Coverage on dataset 1 Coverage on dataset 2
Intervention $\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$ $\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
$a^{\ast}=7x$ 1.00 / 0.19 0.90 / 0.13 0.83 / 0.11 1.00 / 0.02 0.94 / 0.02 0.85 / 0.02
$a^{\ast}=10x$ 1.00 / 0.28 0.91 / 0.23 0.88 / 0.11 1.00 / 0.08 0.84 / 0.07 0.83 / 0.07
Table LABEL:tab:results_unknown presents the empirical coverage of the intervals from our CP method vs. MC dropout across different $\alpha$ and hard interventions $a^{\ast}$. We observe that our conformal prediction intervals are effective and achieve the intended coverage. In contrast, MC dropout does not provide faithful intervals. Our findings are again in line with the literature, where MC dropout is found to produce poor approximations of the posterior and thus might provide poor coverage (Le Folgoc et al., 2021). We present further results in Supplement F.
Figure 6: Prediction intervals for multiple significance levels $\alpha$ for synthetic dataset 1 with intervention $\Delta=5$.
Insights: We plot the prediction intervals across different significance levels $\alpha$ and different covariates $X$ (see Fig. 6). This allows us to inspect the intervals visually. We observe that the intervals behave as expected; that is, the intervals become sharper with increasing significance level $\alpha$. We further see that our CP intervals are slightly wider (see details in Supplement F), yet this is intended because it ensures that the intervals are faithful. For example, our CP intervals (blue) generally include the true outcome. In contrast, the intervals from MC dropout (orange) often do not include the true outcome (e.g., see the bottom row Fig. 6) and are thus not faithful.
5.3 Results for the MIMIC dataset
In Fig. 7, we compare the CP intervals of each two male and female patients of differing ages when treated with increasing duration of mechanical ventilation. Our intervals show higher uncertainty in treatment regions rarely included in the training data (high medium to high treatments). The intervals given by MC-dropout are narrower, which also indicates lower cover- age, thus confirming the effectiveness of our proposed method. This finding aligns with our observation from the synthetic datasets.
Figure 7: CP intervals for potential outcomes of increasing duration of mechanical ventilation for four exemplary patients.
6 Discussion
Limitations: As with any other method, our UQ method has limitations that offer opportunities for future research. Our method relies on the quality of the propensity estimator. Although we incorporate estimation errors in the construction of our intervals, poorly estimated propensities could potentially lead to wide prediction intervals. We acknowledge that our prediction intervals are conservative for the point intervention and for segments of the output space with limited calibration data, implying that a representative calibration dataset is essential for the performance of our method. As for all CP methods, the use of sample splitting may reduce data efficiency.
Broader impact: Our method makes a significant step toward UQ for potential outcomes and, therefore, toward reliable decision-making. To this end, we provided strong theoretical and empirical evidence that our prediction intervals are valid. To this end, our method fills an important demand for using causal ML in medical practice Feuerriegel et al. (2024) and other safety-critical applications with limited data.
Conclusion: We presented a novel conformal prediction method for potential outcomes of continuous treatments with finite-sample guarantees. Our method is applicable to potential outcomes and extends naturally to treatment effects. A key strength of our method is that the intervals are valid under distribution shifts introduced by the treatment assignment, even if the propensity score is unknown and has to be estimated.
Acknowledgements
Appendix A Additional theoretical results
$\gamma_{i}(a):=Y_{i}(a)-Y_{i}(0).$ (21)
Lemma 4.
$C^{+}:=\phi(x_{i},a)+S_{a}^{\ast}-\phi(x_{i},0)+S_{0}^{\ast},$ (22)
$C^{-}:=\phi(x_{i},a)-S_{a}^{\ast}-\phi(x_{i},0)-S_{0}^{\ast}.$ (23)
Proof.
$\varepsilon_{i}(a):=Y_{i}(a)-\phi(x_{i},a).$ (24)
Now observe that
$=$ $P((Y_{i}(a)\geq C^{-}+Y_{i}(0))\land(Y_{i}(a)\leq C^{+}+Y_{i}(0)))$ (27)
Thus, it follows directly that
$P(\gamma_{i}(a)\in C_{\gamma}(X_{i},a))\geq 1-\alpha.$ (31)
In Section 4.2, we formulated the unknown propensity shift in terms of
and minimized over $\sigma$ and $c_{a}$ in Theorem 2 to construct the CP intervals. However, in certain applications, it might be beneficial to control the spread of the approximation of $\delta_{a^{\ast}}(a)$ through fixing $\sigma$ to a small value $\sigma_{0}$ and performing the soft intervention $\tilde{\pi}(a\mid x)=\frac{c_{a}}{\sqrt{2\pi}\sigma_{0}}\frac{\exp{(-\frac{(a-a^{\ast})^{2}}{\sigma_{0}^{2}})}}{\hat{\pi}(a\mid x)}$. In this case, the resulting optimization problem is a convex problem similar to Theorem 1. We present the alternative optimization problem below.
Theorem 3 (Alternative for Theorem 2: Conformal prediction intervals for unknown propensity scores).
[…]
$C(x_{n+1},a^{*}):=\{y\mid S_{n+1}(y)\leq S^{\ast}\}$ (34)
satisfies the desired coverage guarantee
$P(Y(a^{\ast}\in C(X_{n+1},a^{*})\geq 1-\alpha,$ (35)
Proof.
The statement follows from Theorem 1. ∎
Appendix B Algorithm

### v3

@@ 4.1 Scenario 1: Known propensity score
Lemma 4.1 ([18]).
Let $\mathcal{F}$ define a finite-dimensional function class that includes the function $f$ characterizing the shift in the (potentially unknown) propensity function $\pi$ (see Eq. (2)). Define the distribution-shift-calibrated $(1-\alpha)$-quantile of the non-conformity scores as
$\hat{g}_{S}(X_{n+1}):=\arg\min\limits_{g\in\mathcal{F}}\frac{1}{n-m}\left(\sum_{i=m+1}^{n}l_{\alpha}(g(X_{i}),S_{i})+l_{\alpha}(g(X_{n+1}),S)\right)$ (5)
@@
Building upon Lemma 4.1, we derive our first main result in Thm. 4.2. We define the finite-dimensional function class of interest as :=$\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$. It is easy to verify that all $f\in\mathcal{F}$ represent the desired propensity shift to $\tilde{\pi}(a\mid x)=\pi(a+\Delta_{A}\mid x)$ as defined in Eq. 2.
Theorem 4.2 (Conformal prediction intervals for known baseline policy).
@@
$\max_{\eta_{i},i=m+1,\ldots,n+1}\;\;\min_{\theta>0}\ \sum_{i=m+1}^{n}\eta_{i}\,\left(S_{i}-\theta\frac{\pi(a_{i}+\Delta_{A}\mid x_{i})}{\pi(a_{i}\mid x_{i})}\right)+\eta_{n+1}\left(S-\theta\frac{\pi(a^{\ast}\mid x_{n+1})}{\pi(a_{n+1}\mid x_{n+1})}\right)$ (8)
[…]
@@
We provide a full proof in Supplement D.2. Here, we briefly outline the underlying idea of the proof. First, we show that the function class $\mathcal{F}$ indeed satisfies Eq. (2) for the intervention $A^{*}(X)=A+\Delta_{A}$, and we then rewrite Eq. (5) as a convex optimization problem. Next, we exploit the strong duality property. We optimize over the corresponding dual problem to receive a dual prediction set with equal coverage probability. Finally, we derive $S^{\ast}$ from the dual prediction set to construct $C_{n+1}$ and prove the overall coverage guarantee. ∎
@@ 4.2 Scenario 2: Unknown treatment policy
which hinders a direct adaptation of Theorem 4.2 due to the inherent discontinuity of the improper function. Hence, we make the following assumption on the propensity estimator.
@@
can be represented in terms of Eq. (2) through functions $f\in\mathcal{F}$.
@@
See Supplement D.1. ∎
Following the motivation in scenario  1, we thus aim to estimate the $(1-\alpha)$-quantile of the non-conformity scores under the distribution shift in Lemma 4.1. We can reformulate this problem as
@@
s.t. $\quad S_{i}-\frac{c_{a}}{\sqrt{2\pi}\sigma}\frac{\exp{(-\frac{(a_{i}-a^{\ast})^{2}}{2\sigma^{2}})}}{\hat{\pi}(a_{i}\mid x_{i})}-u_{i}+v_{i}=0,\quad\forall i=m+1\ldots,n+1,$
$\quad u_{i},v_{i}\geq 0,\quad\forall i=m+1\ldots,n+1$
for the imputed score $S_{n+1}=S$. As the score is unknown, computing the CP interval would require solving (PS) for all $S\in\mathbb{R}$, yet which is computationally infeasible. As a remedy, we previously exploited properties of the dual optimization problem and the Lagrange multipliers of the convex problem in Theorem 4.2 to efficiently compute the CP intervals (see the proof of Theorem 4.2 in Supplement D). However, the present non-convex problem does not automatically allow for the same simplifications. Instead, we now present a remedy for efficient computation of the CP intervals in the following lemma. We prove Lemma 4.4 in Supplement D.
Lemma 4.4.
The problem (PS) is Type-I invex and satisfies the linear independence constraint qualification (LICQ).
Lemma 4.4 allows us to derive properties of the present non-convex optimization problem in terms of the Karush-Kuhn-Tucker (KKT) conditions. For this, we note that the fulfillment of the LICQ serves as a sufficient regularity condition for the KKT to hold at any (local) optimum of (PS). Combined with the Type-I invexity of the objective function and the constraints, the KKT conditions are not only necessary but also sufficient for a global optimum. As a result, we can employ the KKT conditions at the optimal values [^6] We provide an interpretation of the optimal values in Supplement F. Therein, we further discuss the implications of the proposed kernel smoothing of $\delta_{a^{\ast}}(a)$. $\sigma^{\ast}$ and $c_{a}^{\ast}$ to derive coverage guarantees of our CP interval in a similar fashion as in Theorem 4.2. We thus arrive at the following theorem to provide CP intervals for the scenario with unknown propensity scores.
Theorem 4.5 (Conformal prediction intervals for unknown propensity scores).
@@
s.t. $\quad S_{i}-\frac{c_{a}}{\sqrt{2\pi}\sigma}\frac{\exp{(-\frac{(a_{i}-a^{\ast})^{2}}{2\sigma^{2}})}}{\hat{\pi}(a_{i}\mid x_{i})}-u_{i}+v_{i}=0,\quad\forall i=m+1\ldots,n+1$
$\quad u_{i},v_{i}\geq 0,\quad\forall i=m+1\ldots,n+1$
for an imputed unknown $S_{n+1}=S$. Let $S^{\ast}$ be defined as the maximum $S$ s.t. $v_{n+1}^{S}>0$. Then,
@@
We now use Thm. 4.5 to present an algorithm for computing CP intervals of potential outcomes from continuous treatment variables under unknown propensities in Alg. 1. We present a similar algorithm for scenario  1 with known propensities and discuss the computational complexity in Supplement B.
5 Experiments
Baselines: As we have discussed above, there are no baselines that directly compute prediction intervals with finite-sample guarantees for potential outcomes of continuous treatments. Therefore, we compare our method against MC dropout [14] and deep ensemble methods [33]. Yet, we again emphasize that MC dropout is an ad hoc method with poor approximations of the posterior, which is known to give unfaithful intervals [34]. Furthermore, we report the empirical coverage achieved by intervals from a Gaussian process regression (GP). By doing so, we consider a method that assesses the underlying aleatoric uncertainty.
Implementation: All methods are implemented with $\phi$ as a multi-layer perceptron (MLP) and an MC dropout regularization of rate 0.1. Crucially, we use the identical MLP for both our CP method and MC dropout. Hence, all performance gains must be attributed to the coverage guarantees of our conformal method. In the MC dropout baseline, the uncertainty intervals are computed via Monte Carlo sampling. In scenario 2, we perform conditional density estimation by conditional normalizing flows [51]. Implementation and training details are in Supplement F.
Performance metrics: We evaluate the methods in terms of whether the prediction intervals are faithful [e.g., 21]. That is, we compute whether the empirical coverage of the prediction intervals surpasses the threshold of $1-\alpha$ for different significance levels $\alpha\in\{0.05,0.1,0.2\}$. Additionally, we report the width of the resulting intervals in Supplement G.
5.1 Datasets
Synthetic datasets: We follow common practice and evaluate our methods using synthetic datasets [e.g., 2, 25]. Due to the fundamental problem of causal inference, counterfactual outcomes are never observable in real-world datasets. Synthetic datasets enable us to access counterfactual outcomes and thus to benchmark methods in terms of whether the computed intervals are faithful. Additionally, we perform experiments on the semi-synthic TCGA dataset in Supplement C. We hereby show the applicability of our method to high-dimensional real-world data in a controlled environment.
Medical dataset: We demonstrate the applicability of our CP method to medical datasets on the MIMIC-III dataset [26]. MIMIC-III contains de-identified health records from patients admitted to critical care units at large hospitals. Our goal is to predict patient outcomes in terms of blood pressure when treated with a different duration of mechanical ventilation. We use 8 confounders from medical practice (e.g., respiratory rate, hematocrit). Overall, we consider 14,719 patients, split into train (60%), validation (10%), calibration (20%), and test (10%) sets. Details are in Supplement F.
5.2 Results for synthetic datasets
We consider two synthetic datasets with different propensity scores and outcome functions. Dataset 1 uses a step-wise propensity function and a concave outcome function. Dataset 2 is more complex and uses a Gaussian propensity function and oscillating outcome functions. Both datasets contain a single discrete confounder, a continuous treatment, and a continuous outcome. By choosing low-dimensional datasets, we later render it possible to plot the treatment–response curves so that one can inspect the prediction intervals visually. (We later also show that our method scales to high-dimensional settings as part of the real-world dataset.) Details about the data generation are in Supplement F.
We now evaluate the faithfulness of our CP intervals. On each dataset, we analyze the performance of the prediction intervals in the presence of various soft interventions $\Delta_{a}\in\{1,5,10\}$ and hard interventions $a^{\ast}\in\{7x,10x\}$ for each $X=x$. We average the empirical coverage across 50 runs with random seeds. The results are in Fig. 4 and Fig. 5. Additionally, we report the empirical coverage of the GP.
We make the following observations. First, our CP intervals comply with the targeted significance level $\alpha$ and are therefore faithful. Second, both MC dropout and the deep ensemble method have a considerably lower coverage, implying that the intervals are not faithful. This is in line with the literature, where MC dropout is found to produce poor approximations of the posterior [34]. In particular, the ensemble method is highly unfaithful. Thus, we will not consider this baseline in all the following experiments. Third, our method has only a small
variability in terms of empirical coverage, whereas the empirical coverage of MC dropout varies greatly. This corroborates the robustness of our method. Fourth, the results are consistent for both datasets. Fifth, the GP is only able to capture the true potential outcome in the prediction intervals for small distribution shifts ($\Delta=1$) on dataset 2. However, the empirical coverage is extremely low: $\alpha=0.05$: 0.125; $\alpha=0.1$: 0.125; $\alpha=0.2$: 0.0833. This aligns with our expectations, as the aleatoric uncertainty in our experiments is low. Therefore, the GP intervals are small (average width of 0.1293) and barely valid after the intervention. In sum, this demonstrates our CP method’s effectiveness.
Coverage
Data Intervention $\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
1 $a^{\ast}=7x$ 1.00 / 0.19 0.90 / 0.13 0.83 / 0.11
$a^{\ast}=10x$ 1.00 / 0.28 0.91 / 0.23 0.88 / 0.11
2 $a^{\ast}=7x$ 1.00 / 0.02 0.94 / 0.02 0.85 / 0.02
$a^{\ast}=10x$ 1.00 / 0.08 0.84 / 0.07 0.83 / 0.07
Figure 6: CP intervals for multiple significance levels $\alpha$ for Dataset 1 with intervention $\Delta=5$.
Table 1 presents the empirical coverage of the intervals from our CP method vs. MC dropout across different $\alpha$ and hard interventions $a^{\ast}$. We observe that our CP intervals are effective and achieve the intended coverage. In contrast, MC dropout does not provide faithful intervals. Our findings are again in line with the literature, where MC dropout is found to produce poor approximations of the posterior and thus might provide poor coverage [34]. We present further results in Supplement G.
Insights: We plot the intervals across different levels $\alpha$ and covariates $X$ (Fig. 6), allowing us to inspect the intervals visually. We observe that the intervals behave as expected: they become sharper with increasing significance level $\alpha$. We further see that our CP intervals are slightly wider (see details in Supplement G), yet this is intended as it ensures that the intervals are faithful. Our CP intervals (blue) generally include the true outcome. In contrast, the intervals from MC dropout (orange) often do not include the true outcome (e.g., see the bottom row Fig. 6) and are thus not faithful.
Figure 7: CP intervals for potential outcomes of increasing duration of mechanical ventilation for 4 exemplary patients.
5.3 Results for the MIMIC dataset
In Fig. 7, we compare the CP intervals of two male and two female patients of differing ages when treated with increasing duration of mechanical ventilation. Our intervals show higher uncertainty in treatment regions rarely included in the training data (high medium to high treatments). The intervals given by MC-dropout are narrower, which suggests lower coverage, confirming the effectiveness of our proposed method. This finding aligns with our observation from the synthetic datasets.
6 Discussion
Limitations: As with any other method, our UQ method has limitations that offer opportunities for future research. Our method relies on the quality of the propensity estimator. Although we incorporate estimation errors when constructing intervals, poorly estimated propensities could potentially lead to wide prediction intervals. We acknowledge that our intervals are conservative for point interventions and for segments of the output space with limited calibration data, implying that a representative calibration dataset is essential for the performance of our method. As for all CP methods, the use of sample splitting may reduce data efficiency. Furthermore, we note that the optimization procedure can be computationally expensive for large CATE vectors.
Broader impact: Our method makes a significant step toward UQ for potential outcomes and, thus, toward reliable decision-making. We provided strong theoretical and empirical evidence that our prediction intervals are valid. To this end, our method fills an important demand for using causal ML in medical practice and other safety-critical applications with limited data.
Conclusion: We presented a novel conformal prediction method for potential outcomes of continuous treatments with finite-sample guarantees. Our method extends naturally to treatment effects. A key strength of our method is that the intervals are valid under distribution shifts introduced by the treatment assignment, even if the propensity score is unknown and has to be estimated.
Acknowledgements
Appendix A Additional theoretical results
$\gamma_{i}(a):=Y_{i}(a)-Y_{i}(0).$ (21)
Here, we consider the setting in which the non-conformity score is chosen to be the absolute residual.
Lemma A.1.
$C^{+}:=\phi(x_{i},a)+S_{a}^{\ast}-\phi(x_{i},0)+S_{0}^{\ast},$ (22)
$C^{-}:=\phi(x_{i},a)-S_{a}^{\ast}-\phi(x_{i},0)-S_{0}^{\ast}.$ (23)
Proof.
$\varepsilon_{i}(a):=Y_{i}(a)-\phi(x_{i},a).$ (24)
Now observe that
$=$ $P((Y_{i}(a)\geq C^{-}+Y_{i}(0))\land(Y_{i}(a)\leq C^{+}+Y_{i}(0)))$ (27)
Thus, it follows directly that
$P(\gamma_{i}(a)\in C_{\gamma}(X_{i},a))\geq 1-\alpha.$ (31)
In Section 4.2, we formulated the unknown propensity shift in terms of
and minimized over $\sigma$ and $c_{a}$ in Theorem 4.5 to construct the CP intervals. However, in certain applications, it might be beneficial to control the spread of the approximation of $\delta_{a^{\ast}}(a)$ through fixing $\sigma$ to a small value $\sigma_{0}$ and performing the soft intervention $\tilde{\pi}(a\mid x)=\frac{c_{a}}{\sqrt{2\pi}\sigma_{0}}\frac{\exp{(-\frac{(a-a^{\ast})^{2}}{\sigma_{0}^{2}})}}{\hat{\pi}(a\mid x)}$. In this case, the resulting optimization problem is a convex problem similar to Theorem 4.2. We present the alternative optimization problem below.
Theorem A.2 (Alternative for Theorem 4.5: Conformal prediction intervals for unknown propensity scores).
[…]
$C(x_{n+1},a^{*}):=\{y\mid S_{n+1}(y)\leq S^{\ast}\}$ (34)
satisfies the desired coverage guarantee
$P(Y(a^{\ast})\in C(X_{n+1},a^{*})\geq 1-\alpha,$ (35)
Proof.
The statement follows from Theorem 4.5. ∎
A.3 Soft-interventions on estimated propensities
In the main paper, we presented algorithms for constructing prediction intervals for soft interventions if the propensity function is known and hard interventions if it is unknown. These are arguably the most common scenarios in practice. However, in some cases, one might also be interested in the effect of soft interventions on estimated propensity scores  [e.g., 37]. Therefore, we present an alternative theorem for calculating conformal prediction intervals under soft interventions with estimated propensity scores below.
Theorem A.3 (Conformal prediction intervals for soft interventions with unknown propensity scores).
Let a new datapoint be given with $X_{n+1}=x_{n+1}$ and $A_{n+1}=a_{n+1}$. Furthermore, let $\hat{\pi}$ denote the estimated propensity score with estimation error bounded by $[\frac{1}{M},M]$, $M>0$. The soft intervention is represented by the shift given through $\Delta\in\mathbb{R}$. Let $\eta^{S}=\{\eta_{1}^{S},\ldots,\eta_{n+1}^{S}\}\in\mathbb{R}^{n+1}$ be the optimal solution to
$\max_{\begin{subarray}{c}\eta_{i},\\ i=1,\ldots,n+1\end{subarray}}\min_{\frac{1}{M}\leq c_{a}\leq M}\sum_{i=1}^{n}\eta_{i}\,\left(S_{i}-\frac{c_{a}\hat{\pi}(a_{i}+\Delta\mid x_{i})}{\hat{\pi}(a_{i}\mid x_{i})}\right)+\eta_{n+1}\left(S-\frac{c_{a}\hat{\pi}(a_{n+1}+\Delta\mid x_{n+1})}{\hat{\pi}(a_{n+1}\mid x_{n+1})}\right)$ (36)
[…]
$C(x_{n+1},a^{*}):=\{y\mid S_{n+1}(y)\leq S^{\ast}\}$ (37)
satisfies the desired coverage guarantee
$P(Y(a^{\ast})\in C(X_{n+1},a^{*})\geq 1-\alpha,$ (38)
where with a slight abuse of notation $Y(a^{\ast})$ denotes the potential outcome under the soft intervention represented by $\Delta$.
Proof.
The statement follows from Theorem 4.2 and Theorem 4.5. ∎
Appendix B Algorithm
We now use Thm. 4.5 to present an algorithm for computing CP intervals of potential outcomes from continuous treatment variables under unknown propensities in Alg. 1. We present a similar algorithm for scenario  1 with known propensities and discuss the computational complexity in below.
We make the following comments: In our algorithm, an optimization solver is used to calculate $v_{n+1}$ according to Theorem 4.5. The specific choice of the solver is left to the user. In our experiments in Section 5, we perform the optimization via interior point methods. Further, the overall goal of our algorithm is to find the optimal imputed non-conformity score $S^{\ast}$ such that the coverage guarantees hold. It can be implemented through suitable iterative search algorithms.
@@ 21      $S^{\ast}\leftarrow\frac{S_{\mathrm{up}}+S_{\mathrm{low}}}{2}$;
Algorithm explanation: Theorem 4.5 requires knowledge of the optimal imputed non-conformity score $S^{\ast}$. Since it is unknown beforehand, we implement Algorithm 1 as a binary search algorithm. We find suitable upper and lower bounds for $S^{\ast}$ (until line 8). To validate that the bounds are indeed smaller/larger than the optimal non-conformity score, we observe the corresponding dual value $\nu_{n+1}$ (Theorem 4.5). After finding valid bounds, we start searching for $S^{\ast}$ via standard binary search, continuously increasing the lower and decreasing the upper bound. When the difference between the bounds is less than a specified error tolerance $\varepsilon$, we stop and take $S^{\ast}$ to be the mean of the interval. In every iteration (when calling ‘solver’), we make use of the optimization in Theorem 4.5 and check whether the dual value fulfills the requirement in the Theorem.